OpenStudy (anonymous):

MEDAL!! What is -2 & 3/8 expressed as an improper fraction? A. B. C. D.

jimthompson5910 (jim_thompson5910):

ignore the negative for now

jimthompson5910 (jim_thompson5910):

2 & 3/8 = 2 + 3/8 2 & 3/8 = 2*(8/8) + 3/8 2 & 3/8 = 16/8 + 3/8 2 & 3/8 = ???

OpenStudy (anonymous):

i dont get it? :( @jim_thompson5910

jimthompson5910 (jim_thompson5910):

which step are you stuck on?

OpenStudy (anonymous):

i dont get the way u explained it @jim_thompson5910

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):